# QuEST: Low-bit Diffusion Model Quantization via Efficient Selective Finetuning

### v2

@@ 3.3 Quantization via Efficient Selective Finetuning
In this section, we first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning. Then we introduce a finetuning method QuEST for diffusion models that can significantly boost low-bit performance with less time and memory usage. Figure 1 illustrates the proposed framework, which is a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization. To further address the loss of temporal information due to quantization, we propose an efficient time-aware activation quantizer.
@@ 3.3.1 Sufficiency of Finetuning
We first review the underlying theoretical foundation underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach. Denote the diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the task loss as $L(z_{n,t},\overline{z}_{n,t})$, where $n$ is the number of layers and $\overline{z}_{n,t}$ is the ground-truth. $L$ can be any loss function and here we use the mean squared error (MSE). We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
$\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (4)
where $\Delta$ is the activation perturbation, $\overline{\mathbf{g}}^{(\mathbf{z})}$ is the gradient and $\overline{\mathbf{H}}^{(z_{n,t})}$ is the Hessian matrix. According to (Li et al., 2021; Yuan et al., 2022), for a well-trained model, $\overline{\mathbf{g}}^{(z_{n,t})}=\nabla_{z_{n,t}}L$ approaches 00. Thus the above equation can be simplified to:
@@
This leads to a rather straightforward conclusion: Quantization error can be reduced by reconstructing the model output. However, under low-bit settings, the reasoning from Equation 3.3.1 to Equation 5 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion. Thus we have the following proposition:
@@
Theorem 3.2 indicates that: To minimize the quantization error, ideally we can finetune $\mathbf{w}_{n}$ so that for any $i$, the weights can fit to the data/activation $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$ which the full-precision counterpart may not fit, making the model more robust to varying inputs. In other words, we are finetuning the model weights for better robustness towards large input activation perturbations, therefore enabling easier quantization. We also find that the second term in Equation 3.2 can be minimized simultaneously with the first term, so that the total quantization error can be reduced via a common goal: Activation output alignment via finetuning $\mathbf{w}_{n}$.
Figure 3: Effect of decreasing different activations’ bit-width on the model performance in Stable Diffusion. FeedForward layers experience generation failures at 6 bits, while all other linear layers barely fail at 4 bits and all convolution layers only fail at 4 bits.
3.3.2 Quantization via Selective Finetuning
Data-Free Partial-network-wise Training. To alleviate the need of substantial training data, we propose to adopt the general setting of Post-Training Quantization and construct the calibration set in a data-free manner. By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps using Equation 1, we can obtain the calibration data needed for finetuning the quantized model. In practice, the number of samples generated for each time step is 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
Moreover, inspired by the fact that we can view the activation quantization parameters as additional model parameters which all contribute to the final output, we propose a partial-network-wise training strategy. Different from quantization methods using layer-wise or block-wise reconstruction (Shang et al., 2023; Li et al., 2023a) that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step and only train partial weight parameters (e.g. 7% of the parameters in LDM-4). As depicted in Figure 1, the majority of the model weights remain unchanged at time step $t$, and only a small subset of parameters (including layer weights and scaling factors) related to this time step are updated using the full-precision model’s outputs as ground-truth. The choices for the weights to be finetuned are discussed in the following sections. Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, our approach update the required parameters at once. In this way, we significantly save the time and memory needed for quantization. Note that the quantization parameters for the weights are not updated once initialized, in order to maintain memory efficiency during deployment.
Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. ”N/A” denotes generation failure. ”FP” denotes the full precision model.
Dataset Method
Bit-width
(W/A)
Size
@@ 3.3.2 Quantization via Selective Finetuning
LSUN- Bedrooms (LDM-4) FP 32/32 1045.6 2.95
PTQ4DM 8/8 279.1 4.75
Q-Diffusion 8/8 279.1 4.53
PTQ-D 8/8 279.1 3.75
Ours 8/8 279.1 3.07
PTQ4DM 4/8 148.4 -
Q-Diffusion 4/8 148.4 5.37
PTQ-D 4/8 148.4 5.94
Ours 4/8 148.4 3.26
PTQ4DM 4/4 148.4 N/A
Q-Diffusion 4/4 148.4 N/A
PTQ-D 4/4 148.4 N/A
Ours 4/4 148.4 5.64
LSUN- Churches (LDM-8) FP 32/32 1125.4 4.02
PTQ4DM 8/8 330.6 5.54
Q-Diffusion 8/8 330.6 6.94
PTQ-D 8/8 330.6 6.40
Ours 8/8 330.6 6.55
PTQ4DM 4/8 189.9 -
Q-Diffusion 4/8 189.9 7.80
PTQ-D 4/8 189.9 7.33
Ours 4/8 189.9 7.33
PTQ4DM 4/4 189.9 N/A
Q-Diffusion 4/4 189.9 N/A
PTQ-D 4/4 189.9 N/A
Ours 4/4 189.9 11.76
Efficient Time-aware Activation Quantization. Temporal information is a unique and essential factor in diffusion models. As highlighted in (Shang et al., 2023; Li et al., 2023a), the variation of activation distribution over different time steps is fatal for conventional PTQ methods. Also, as discussed in Section 3.2, accurate time embedding quantization is critical. To discriminate between different time steps, we design the Time-aware Activation Quantizer (TAQuant) that adopts different sets of activation quantization parameters for different time steps. While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost. Inspired by the fact that neighboring time steps yield similar functions in diffusion models (Choi et al., 2022), we use the simple strategy of clustering neighboring time steps so that they share the same calibration data and quantization parameters. Then the set of activation quantization parameters are formulated as:
$\mathbf{s}=\{\mathbf{s}^{t}(x_{t}),\mathbf{s}^{2t}(x_{2t}),...,\mathbf{s}^{T}(x_{T})\},$ (7)
where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $T/t$ adjacent ones. Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance. Section 4.3 further provides an analysis on the trade-off between time efficiency and model performance.
Table 4: Class-conditional image generation results of the LDM-4 model on ImageNet 256$\times$256, containing 1,000 categories.
Bit-width
(W/A)
Method
Size
@@
32/32 FP 1529.7 11.28 7.70 364.73
8/8 Q-Diffusion 428.7 10.60 9.29 350.93
PTQ-D 428.7 10.05 9.01 359.78
Ours 428.7 10.43 6.07 365.12
4/8 Q-Diffusion 237.5 9.29 9.29 336.80
PTQ-D 237.5 8.74 7.98 344.72
Ours 237.5 8.48 6.55 346.02
4/4 Q-Diffusion 237.5 N/A N/A N/A
PTQ-D 237.5 N/A N/A N/A
Ours 237.5 5.98 7.93 202.45
Table 5: Text-to-image generation results (512$\times$512) of Stable Diffusion v1.4 using COCO2014 validation set prompts.
[…]
32/32 FP 3279.1 31.50
8/8 Q-Diffusion 949.0 31.43
Ours 949.0 31.47
4/8 Q-Diffusion 539.1 31.39
Ours 539.1 31.50
4/4 Q-Diffusion 539.1 N/A
Ours 539.1 28.85
Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 in LSUN-Bedrooms. Our method not only succeed well in generating low-bit images, but also aligns more closely with the full-precision counterpart.
During a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations. This implies that the time information operates independently from the primary network flow. Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}^{t}$ using:
$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}^{t})||^{2}],$ (8)
where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$, $\overline{O}(e(t);\mathbf{w}_{l})$ is the activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}^{t})$ is the quantized output. This loss function indicates that the chosen weight parameters are consistently updated across different time steps to ensure robustness to diverse inputs, whereas only the subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
Apart from time embedding layers, attention-related layers are also vital. As depicted in Figure 3, reducing activation bit-width of these layers degrades model performance the most dramatically. Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize $\mathbf{w}_{l}$ and $\mathbf{s}^{t}$ but exclude the quantization parameters already optimized in Equation 8:
$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}^{t}\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (9)
Furthermore, we aim not only for each independent layer’s quantization error reduction, but also for the final generated images’ quality improvement. Therefore, the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced, where the ground-truth is approximated by the output of the full-precision counterpart. By integrating Equation 8 and Equation 9, the final objective is formulated as follows:
$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}^{t})||^{2}]).$ (10)
4 Experiments
4.1 Experiment Settings
To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets (Yu et al., 2015), class-conditional image generation on ImageNet (Deng et al., 2009), and text-to-image generation. The model architectures we quantize include LDM and Stable Diffusion (Rombach et al., 2021), and uses ”WnAm” to represent the quantization setting: n-bit weight quantization and m-bit activation quantization. DDIM samplers (Ho et al., 2020) are adopted for LDMs and the PLMS sampler (Liu et al., 2022) is used for Stable Diffusion. We generate 256 samples per time step for constructing the calibration set, and train 20 epochs over each time step. The Adam optimizer (Kingma & Ba, 2017) is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
[…]
4.2 Experiment Results
Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively. Results are shown in table 3. We compare with methods including PTQ4DM (Shang et al., 2023), Q-Diffusion (Li et al., 2023a) and PTQ-D (He et al., 2023b). The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet (Li et al., 2023a).
Table 6: Different component comparison and efficiency comparison on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
@@ 4.2 Experiment Results
PTQ (Li et al., 2023a) 4/4 5120 23.08 10334 148.4 428.8
TAQuant 4/4 5120 11.52 9822 148.4 12.56
+ $\text{SLA}_{\text{TE}}$ 4/4 5120 13.13 12134 148.4 6.41
+ $\text{SLA}_{\text{A}}$ 4/4 5120 15.25 12450 148.4 5.78
Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps. Table 5 shows the results. Images are generated based on the 10,000 prompts sampled from the COCO2014 (Lin et al., 2015) validation set, and CLIP Score is calculated based on the ViT-B/16 backbone. Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
4.3 Experimental Analysis
Efficiency comparison and effects of different component. In Table 6, we present a comparison of efficiency between our method and the post-training quantization approach on the LSUN-Bedrooms dataset. Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in memory usage. We also show how each proposed component improves the generation performance. The results indicate that sequentially finetuning time embedding layers followed by attention-related layers leads to consistent improvements in model performance.
Table 7 presents the comparison of with and without the use of task loss $\mathcal{L}_{d}$. The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively. However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.
Table 7: Influence of task loss supervision on model performance.
Method
$\text{SLA}_{\text{TE}}$
w/o $\mathcal{L}_{d}$
$\text{SLA}_{\text{TE}}$
w/ $\mathcal{L}_{d}$
$\text{SLA}_{\text{A}}$
w/o $\mathcal{L}_{d}$
$\text{SLA}_{\text{A}}$
w/ $\mathcal{L}_{d}$
[…]
Time step sampling frequency analysis. While time steps are sampled uniformly, will the performance consistently improve when we sample more frequently? Table 8 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance. Metrics are calculated using 10,000 samples generated from LDM-4 on LSUN-Bedrooms. The results show that the though performance improves consistently, the amount of growth gradually decreases. Starting from 5 time steps, using 15 more time steps leads to 25.22 sFID improvement. But using another 30 time steps (from 20 to 50) only leads to 2.92 sFID improvement. Given the linearly increasing time cost associated with using more time steps, we opt for only 20 out of 200 time steps to enhance quantization efficiency.
Table 8: Effect of number of uniformly sampled time steps for W4A4 LDM-4 on LSUN-Bedrooms dataset.
T-Num 5 10 20 25 40 50
FID $\downarrow$ 19.34 15.27 12.56 12.16 11.04 10.54
sFID $\downarrow$ 54.00 37.47 28.78 27.99 26.42 25.86

### v3

@@ 2.3 Quantization via Efficient Selective Finetuning
We first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning. Then we introduce a finetuning method QuEST for diffusion models that can significantly boost low-bit performance with less time and memory usage. Figure 1(c) shows our framework, being a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization. To further address the loss of temporal information due to quantization, an efficient time-aware activation quantizer is proposed.
@@ 2.3.1 Sufficiency of Finetuning
Motivated by the phenomenon that current PTQ methods completely fail when applied to low-bit diffusion models and driven by the phenomenon that the performance degradation caused by activation bit-width decrease is much more significant than weight bit-width (8-bit to 4-bit) in diffusion models, we would like to explore theoretically why previous PTQ methods fail and how to overcome this.
We first review the underlying theory underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach. Denote the diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the task loss as $L(z_{n,t},\overline{z}_{n,t})$, where $n$ is the number of layers and $\overline{z}_{n,t}$ is the ground-truth. $L$ can be any loss function and here we use the mean squared error (MSE). We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (3)
where $\Delta$ is the activation perturbation, $\overline{\mathbf{g}}^{(\mathbf{z})}$ is the gradient and $\overline{\mathbf{H}}^{(z_{n,t})}$ is the Hessian matrix. According to [22, 40], for a well-trained model, $\overline{\mathbf{g}}^{(z_{n,t})}=\nabla_{z_{n,t}}L$ approaches 00. Thus the above equation can be simplified to:
@@
This leads to a rather straightforward conclusion: Quantization error can be reduced by reconstructing the model output. However, under low-bit settings, the reasoning from Equation 3 to Equation 4 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion. Thus we have the following proposition:
@@
$\approx$ $2\epsilon^{T}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{FP}})+\frac{1}{2}\sum_{i=1}^{K}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})^{T}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})$ (5)
Proof is provided in Appendix B. Theorem 2.2 indicates that: To minimize the quantization error, ideally we can finetune $\mathbf{w}_{n}$ so that for any $i$, the weights can fit to the data/activation $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$ which the full-precision counterpart may not fit, making the model more robust to varying inputs. In other words, we are finetuning the model weights for better robustness towards large input activation perturbations, therefore enabling easier quantization. We also find that the second term in Equation 2.2 can be minimized simultaneously with the first term, so that the total quantization error can be reduced via a common goal: Activation output alignment via finetuning $\mathbf{w}_{n}$.
2.3.2 Quantization via Selective Finetuning
Data-Free Partial-network-wise Training. To alleviate the need of substantial training data, we construct the calibration set in a data-free manner. By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps, we can obtain the calibration data needed for finetuning the quantized model. In practice, the number of samples generated for each time step is 128 or 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
As depicted in Figure 1(c), to overcome the quantization difficulty demonstrated in property efficiently, we update partial model weights ($\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$) that only account for a small subset of parameters related to the time step $t$. The other weight parameters (such as $\mathbf{W}_{l^{{}^{\prime}}}$, $\mathbf{W}_{l^{{}^{\prime\prime}}}$) are frozen during optimization. For example on LDM-4 [32], at most 7% of the parameters are included for adjustment. The choices for the weights to be finetuned are discussed in the following sections. To better conserve the performance and behavior of the full-precision model, we use both its intermediate activations and final outputs as the ground-truth for supervision. Figure 4 visualizes the generation results and demonstrates the effectiveness of this strategy in preserving image details.
The activation quantization parameters can be viewed as additional model parameters, therefore we further propose a network-wise training strategy. Different from quantization methods using layer-wise or block-wise reconstruction [19, 34] that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step together with the partial weight parameters. Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, we update the required parameters at once. In this way, we significantly save the time and memory needed for quantization. Note that the quantization parameters for the weights are not updated once initialized, in order to maintain memory efficiency during deployment.
Efficient Time-aware Activation Quantization. As discussed in property , accurate time embedding quantization is critical. To discriminate between different time steps, we design a temporal quantizer specifically for activation quantization that adopts different sets of quantization parameters for different time steps. While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost.
Motivated by the fact that neighboring time steps yield similar functions in diffusion models [3], we cluster neighboring time steps so that they share the same quantization parameters and are calculated by the same calibration data $x_{t}$: $\mathbf{s}(t)=\mathbf{s}(t+1)=...=\mathbf{s}(t+\Delta t)$, where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $\Delta t$ adjacent ones. Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance. Section 3.2 further provides an analysis on the performance and efficiency trade-off.
Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. "N/A" denotes generation failure. "FP" denotes the full precision model.
[…]
Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 for LSUN-Bedrooms. Our method not only succeed well in generating low-bit images, but also aligns more closely with the full-precision counterpart.
In a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations. This implies that the time information operates independently from the primary network flow. Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}(t)$:
$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))||^{2}],$ (6)
where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$. $\overline{O}(e(t);\mathbf{w}_{l})$ is the final activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))$ is the quantized output. This objective function indicates that the chosen weight parameters are consistently updated across different time steps, so as to ensure robustness to diverse temporal inputs. Nevertheless, only the small subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
Apart from time embedding layers, attention-related layers are also vital (according to property ). As depicted in Figure 3, reducing the activation bit-width of these layers degrades model performance the most dramatically. Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize the corresponding weights $\mathbf{w}_{l}$ and quantization parameters $\mathbf{s}(t)$. However, we exclude the quantization parameters already optimized in Equation 6:
$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}(t)\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (7)
Furthermore, we aim not only for each independent layer’s quantization error reduction, but also for the final generated images’ quality improvement. Therefore, we adopt the Task Loss Guidance strategy where the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced and the ground-truth is approximated by the output of the full-precision counterpart. By integrating Equation 6 and Equation 7, the final objective is formulated as follows:
$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}(t))||^{2}]).$ (8)
3 Experiments
3.1 Experiment Settings
To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets [39], class-conditional image generation on ImageNet [5], and text-to-image generation. The model architectures we quantize include LDM and Stable Diffusion [32], and uses "WnAm" to represent the quantization setting: n-bit weight quantization and m-bit activation quantization. DDIM samplers [15] are adopted for LDMs and the PLMS sampler [26] is used for Stable Diffusion. We generate 256 samples per time step for constructing the calibration set, and train 20 epochs over each time step. The Adam optimizer [18] is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
Table 4: Class-conditional generation results using LDM-4 on ImageNet 256$\times$256.
@@ 3.1 Experiment Settings
3.2 Experiment Results and Analysis
Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively. Results are shown in table 3. The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet [19].
Table 5: Text-to-image generation on Stable Diffusion (512$\times$512) using COCO2014 prompts.
@@ 3.2 Experiment Results and Analysis
Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps. Table 5 shows the results. Images are generated based on the 10,000 prompts sampled from the COCO2014 [24] validation set, and CLIP Score is calculated based on the ViT-B/16 backbone. Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
Table 6: Influence of task loss supervision on model performance.
@@
Efficiency comparison and effects of components. Table 6 compares with and without the use of task loss $\mathcal{L}_{d}$. The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively. However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.
Table 7 compares efficiency and performance between our method and the post-training quantization approach on the LSUN-Bedrooms dataset when using LDM-4. Our approach, while utilizing the same quantity of calibration data as the PTQ method, is more time-efficient and incurs only a 20% increase in GPU memory usage. We also show how each proposed component improves the generation performance. The results indicate that sequentially finetuning time embedding layers followed by finetuning attention-related layers leads to consistent performance improvements.
Table 7: Component and efficiency comparisons on LDM-4 (LSUN-Bedrooms 256 $\times$ 256).
@@
PTQ [19] 4/8 5120 23.08 10334 148.4 5.37
Baseline 4/8 5120 11.52 9822 148.4 6.95
+ $\text{SLA}_{\text{A}}$ 4/8 5120 13.13 12134 148.4 4.41
+ $\text{SLA}_{\text{A}}$ + $\text{SLA}_{\text{TE}}$ 4/8 5120 15.25 12450 148.4 3.26
Figure 5: Effect of uniformly sampled time steps number. Circle size represents quantization computation cost. Red square indicates the value we choose.
Time step sampling frequency analysis. While time steps are sampled uniformly, will the performance consistently improve when we sample more frequently? Figure 5 illustrates the trade-off between the increased number of sampled time steps (’T-Num’) and generation performance. Metrics are calculated using 10,000 samples generated from LDM-4 on LSUN-Bedrooms. The results show that the though performance improves consistently, the amount of growth gradually decreases. Starting from 5 time steps, using 15 more time steps leads to 25.22 sFID improvement. But using another 30 time steps (from 20 to 50) only leads to 2.92 sFID improvement. Given the linearly increasing time cost associated with using more time steps, we opt for only 20 out of 200 time steps to enhance quantization efficiency.
@@
Integration with LoRA finetuning. Finetuning quantized models is not trivial. Thus, we use QALoRA [11] that extends LoRA [16] and finetune on the ImageNet 256$\times$256 dataset. The weight rank of 32 is used for the LoRA modules and the parameters are trained over 100 time steps for 160 epochs. We find that integrating the QALoRA technique leads to a 5.62 decrease in FID, indicating that finetuning the original layers is a better solution for performance preservation.